Plastic division has an expansion opportunity that costs 550,000 with a residual value of 10,000 and a useful life of 5 years. Assume Straight Line depreciation for accrual purposes and MACRS for taxes. The cost of capital (minimum required return) is 13 percent and the effective tax rate is 29 percent. The annual sales increase due to this project is 320,000 with 90 percent in cash for the year and a contribution margin ratio of 50 percent. The increase in working capital is 9,000 dollars, which will be fully returned at the end of the useful life of the project. Their marginal tax rate is 21 percent. Find the Accrual Accounting Rate of Return for the project WITHOUT the effect of income taxes.
Answer #1

Answer:

Accrual purpose: Depreciation method is Straight Line method

Annual depreciation = (Cost - Salvage value) / Useful life = (550000 - 10000) / 5 = $108,000

Annual Contribution margin = Accrual Sales * contribution margin ratio = 320000 * 50% = $160,000

Annual Income before tax = Annual Contribution margin - Annual depreciation = 160000 - 108000 = $52,000

Annual income before tax remains same for all 5 years; hence

Average Annual Income before tax = $52,000

Average Investment = (550000 + 10000) / 2 = $280,000

Accrual Accounting Rate of Return for the project WITHOUT the effect of income taxes =Average Annual Income before tax / Average Investment

= 52000 / 280000

= 18.57%

Accrual Accounting Rate of Return for the project WITHOUT the effect of income taxes = 18.57%